OpenStudy (anonymous):

Integral of (y^2+4) * y^-1 dy. I think this by parts.

myininaya (myininaya):

multiply (y^2+4) and y^(-1) then integrate

myininaya (myininaya):

\[\int\limits_{}^{}(y+4y^{-1}) dy=?\]

OpenStudy (anonymous):

(Y^2+4)/y

OpenStudy (zale101):

\[\int\limits_{}^{}y~dy+\int\limits_{}^{}\frac{4}{y}~dy\] Integration by parts is not needed, it's just linearity